OpenStudy (anonymous):

given that f(x)=-9x+3 and g(x)=x^4, find (f.g)(x) and its domain

OpenStudy (anonymous):

Just as clarification of notation: Does: \[(f.g)(x)=f(g(x))\]

OpenStudy (anonymous):

no its like (fog)(x)

OpenStudy (anonymous):

its like that but i cant get it on the keyboard

OpenStudy (anonymous):

help mepls

OpenStudy (anonymous):

Ahhh, I see, you are looking for f of g of x. This can be written as below. \[f(g(x))\] We can also write this as f(x) where we let x = g(x): \[f(x)\] So substitute in x = g(x) \[f(g(x))\] Now we'll try this working with your functions. So we have: \[f(x)=-9x+3\] Now we need to let x = g(x) So: \[x=x^4\] Now we substitute this new value for x in to the function for f(x) to give our final answer: \[f(g(x))=-9x^4+3\] Helpful?

OpenStudy (anonymous):

not really can you give me the answer

OpenStudy (anonymous):

Okay, you think you can substitute the value of g(x) into f(x) as if it was just an x value. That will give you the answer and if you look at my working above I have done the answer. Is there anything in particular that you don't understand about my explanation? :)

OpenStudy (anonymous):

oh so you just put x^4 into the 9x right?

OpenStudy (anonymous):

Yeah, you substitute it in as if it was an x value. It can get quite a bit harder than that though :)

OpenStudy (anonymous):

whats the domain then?

OpenStudy (anonymous):

The domain would be real numbers because it's just a quartic function

OpenStudy (anonymous):

infinity, negative infinity?

OpenStudy (anonymous):

Yep, when you think about the function there is no number you can give it that won't return a real answer back.

OpenStudy (anonymous):

i have so many more questions

OpenStudy (anonymous):

You can close this question and open another one if you would like to ask another question. I'll keep a look out for it :)